OpenStudy (anonymous):

Is my answer correct??? Solve. x2 + 8x = 24 -4 _+_ 2 sq root 5

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

\[x^2+8x=24\] \[(x+4)^2=24+16=40\] \[x+4=\pm\sqrt{40}=\pm2\sqrt{10}\] \[x=-4\pm2\sqrt{10}\]

OpenStudy (anonymous):

I got 10, my error!

OpenStudy (anonymous):

whew!

OpenStudy (anonymous):

At fist I got sq root 5, but i did a step wrong!!!! ;( and got 10

OpenStudy (anonymous):

first**

OpenStudy (anonymous):

got it now right?

OpenStudy (anonymous):

yes sir! thank you <3

OpenStudy (anonymous):

yw (sir.., i like that...)

OpenStudy (anonymous):

-4 _+_ sq root 10! :)

OpenStudy (anonymous):

you're welcome, the least I can do, sir.....! ;)

OpenStudy (anonymous):

\(-4\pm2\sqrt{10}\)

OpenStudy (anonymous):

tehehehe :)))

OpenStudy (anonymous):

don't forget the 2 in front of the square root

OpenStudy (anonymous):

I won't ;)